M&M's sell for $6 a pound. Skittles are $3 a pound. How many pounds of M&M's should be mixed with 12 pounds of skittles to get a mixture selling for $4.20 a pound?

Answer by checkley79(3341)

6M+3*12=4.20(12+M)
6M+36=50.4+4.20M
6M-4.2M=50.4-36
1.8M=14.4
M=14.4/1.8
M=8 POUNDS OF M&M IS USED.
PROOF:
6*8+3*12=4.20(12+8)
48+36=4.20*20
84=84